HUBBLE SPACE TELESCOPE DAILY REPORT #1995 PERIOD COVERED: 0000Z (UTC) 10/30/97 - 0000Z (UTC) 10/31/97 Daily Status Report as of 304/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of NIC/1/2/3 7596 (NICMOS MULTIACCUM Dark Characterization) The Near-Infrared Camera Multi-Object Spectrograph (Cameras One, Two and Three) was used to make a NICMOS MULTIACCUM dark characterization. The purpose of this proposal is to monitor the darks using three different MULTIACCUM sequences. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Seven Sets of STIS/CCD 7601 (CCD Dark and Bias Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to conduct bias and dark monitoring. Also, observations were made to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were completed as planned, and no anomalies were reported. 1.4 Completed NIC/1/2/3 7447 (The Cosmological Evolution of Quasar Host Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galaxy MC2112+172 as part of a survey to study the cosmological evolution of galaxies which host quasars. Other point spread function observations were taken to test the calibration of the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 7782 (STIS Non-Scripted Parallel Archive Proposal - Half Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one-half orbit). The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 1.7 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of NIC/1/2 7243 (High Resolution Imaging of Io) The Near Infrared Camera (Cameras One and Two) was used to make observations of Jupiter's moon Io. The technique of super-resolution imaging will be used to obtain near-IR images of Io with sufficient resolution to map the surface on this volcanically active satellite of Jupiter. Observations were also made of standard star HD201592. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/2 7495 (Survey of Gravitational Lenses as Cosmological Tools) The Near Infrared Camera (Camera Two) was used to make survey observations of object MG2016+112 by obtaining deep, high-resolution images in the optical and near IR in order to see if it is gravitationally lensed and could be used as a cosmological tool. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Four Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.11 Completed WF/PC-2/NIC/3 7563 (Francis Cluster) The WF/PC-2 and the Near Infrared Camera (Camera Three) was used to make observations of the cluster Gal-Clus-2 (a.k.a., Francis Cluster) using various modes. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.12 Completed WF/PC-2 7431 (Snapshot Survey of Microlensed Source Stars) The WF/PC-2 was used to make snapshot survey images of microlensed source star Macho-96-B. Gravitational microlensing surveys have revealed an excess of lensing events over theoretical predictions toward both the Large Magellanic Cloud and the Galactic bulge. The observations were completed as planned, and no anomalies were reported. 1.13 Completed FGS 6479 (Masses of Low Luminosity Hyades Cluster Membe7782 The FGSs were used to make astrometric observations of low luminosity star VA351-5 which is a member to the Hyades Cluster. This cluster, owing to its nearness, richness, and young age, occupies a pivotal role in the studies of star formation and stellar populations. The observations were executed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 4 Per HSTAR 7227, the acquisition at 304/054436Z, which occurred in a ZOE, was attempting to achieve/maintain fine lock (FL) at AOS. At astrometry start time, the FL bits of both FGS1 and FGS2 were set, both stop bits were set, and the FGSs were not under SSM control. All bits remained constant during astrometry, but upon its conclusion, the FGS loss of lock flags flagged and SSM control was eventually entered at 304/063409Z, possibly affecting FGS and NICMOS proposals (which will be detailed in the next report). Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: To support a possible ARU/PRT uplink connected with the acquisition failure detailed in 2.1, HST operations personnel performed an operations request which switched communications to the LGAs at 304/0640Z. When it was determined that the procedure was not required, communications were returned to the HGAs at 30/0703Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH